Media Services charges $20 for a phone and $30/month for its economy plan. The equation c=30t+20 describes the total cost, c, of operating a Media Services phone for t months.
The total cost of services for three months is __.
If a customer only has $90 avaiable, how many months of service can she recieve?
a) 2 1/2
b) 2 1/3
c) 2
d) 3

Answer by stanbon(75887)

Media Services charges $20 for a phone and $30/month for its economy plan. The equation c=30t+20 describes the total cost, c, of operating a Media Services phone for t months.
The total cost of services for three months is
c(3) = 30(3)+20 = 90+20 = 110.
=========================
If a customer only has $90 avaiable, how many months of service can she recieve?
Solve: 90 = 30t+20 for "t":
70 = 30t
t = 7/3 = 2 1/3 months
===============================
Cheers,
Stan H.
===============================

a) 2 1/2
b) 2 1/3
c) 2
d) 3